# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ Abstract
While conditional diffusion models have achieved remarkable success in various applications, they require abundant data to train from scratch, which is often infeasible in practice. To address this issue, transfer learning has emerged as an essential paradigm in small data regimes. Despite its empirical success, the theoretical underpinnings of transfer learning conditional diffusion models remain unexplored. In this paper, we take the first step towards understanding the sample efficiency of transfer learning conditional diffusion models through the lens of representation learning. Inspired by practical training procedures, we assume that there exists a low-dimensional representation of conditions shared across all tasks. Our analysis shows that with a well-learned representation from source tasks, the sample complexity of target tasks can be reduced substantially. In addition, we investigate the practical implications of our theoretical results in several real-world applications of conditional diffusion models. Numerical experiments are also conducted to verify our results.
@@ 1 Introduction
Conditional diffusion models (CDMs) utilize a user-defined condition to guide the generative process of diffusion models (DMs) to sample from the desired conditional distribution. In recent years, CDMs have achieved groundbreaking success in various generative tasks, including text-to-image generation [Ho et al., 2020; Song et al., 2020; Ho and Salimans, 2022; Rombach et al., 2022], reinforcement learning [Janner et al., 2022; Chi et al., 2023; Wang et al., 2022; Reuss et al., 2023], time series [Tashiro et al., 2021; Rasul et al., 2021], and life science [Song et al., 2021; Watson et al., 2022; Gruver et al., 2024; Guo et al., 2024].
Training a CDM from scratch requires a large amount of data to achieve good generalization. However, in practical scenarios, users often have access to only limited data for the target distribution due to cost or risk concerns, making the model prone to over-fitting. In such small data regime, transfer learning has emerged as a predominant paradigm [Moon et al., 2022; Ruiz et al., 2023; Xie et al., 2023; Han et al., 2023]. By leveraging knowledge acquired during pre-training on large source datasets, transfer learning enhances the performance of fine-tuning on target tasks, facilitating few-shot learning and significantly improving practicality.
Among the successful applications of transfer learning CDMs, the conditions are typically high-dimensional vectors with embedded low-dimensional representations (features) that encapsulate all the information required for inference. In addition, these representations are likely to be task-agnostic, enabling effective knowledge transfer. For example, in text-to-image generation, the text input is inherently in high-dimensional space, but contains low-dimensional semantic information such as object attributes, spatial relationships, despite the differences of styles or contents in different image distributions. To take advantage of this structure, condition encoders are often frozen in the fine-tuning stage [Rombach et al., 2022; Esser et al., 2024], which typically constitutes a significant portion of the overall model (see Table 1).
@@
Text-to-Image [Esser et al., 2024] 2-8B 4.7B
Text-to-Audio [Liu et al., 2024] 350-750M 750M
Reinforcement Learning [Chi et al., 2023] 9M 20-45M
@@
There are some recent works attempting to study the theoretical underpinnings of CDMs [Fu et al., 2024; Jiao et al., 2024; Hu et al., 2024], but focus on single task training. Notably, Yang et al. [2024] investigates transfer learning DMs under the assumption that the data is a linear transformation of a low-dimensional latent variable following the same distribution across all tasks. However, fine-tuning merely the data encoder is not a widely adopted training approach in practice.
@@
  • In Section 3, we establish the first generalization guarantee for transferring score matching error in CDMs, showing that transfer learning can reduce the sample complexity for learning condition encoder in the target task. This is aligned with existing transfer learning theory in supervised learning. Specifically, we present two results in Theorem 3.4 and Theorem 3.6, under the settings of task diversity assumption and meta-learning [^1] In practice, the terms such as transfer learning, meta-learning, learning-to-learn, etc., often refer to the same training paradigm, i.e., to fine-tune on target tasks with limited data using knowledge from source tasks. However, in the theoretical framework, we use the term meta-learning to emphasize that target tasks and source tasks are randomly sampled from a meta distribution [Baxter, 2000], whereas in transfer learning, the tasks are fixed., respectively. On the technical side, we develop a novel approach to tackle Lipschitz continuity under weaker assumptions on data distribution in Lemma 3.1, which may be of independent interest for the analysis of even single-task diffusion models.
@@
  • In Section 5, we further utilize our results to establish statistical guarantees in practical applications of CDMs. In particular, we investigate amortized variational inference (Theorem 5.1) and behavior cloning (Theorem 5.2), and present guarantees in terms of posterior estimation and optimality gap, laying the theoretical foundations of transfer learning CDMs in practice. We also conduct numerical experiments in Section 6 to verify our results.
@@ 1.1 Related Works
Recently, some works analyze the score approximation theory via deep neural networks and corresponding sample complexity bounds for diffusion models. Oko et al. [2023] considers distributions with density in Besov space and supported on bounded domain. Chen et al. [2023b] assumes the data distribution lies in a low-dimensional linear subspace and obtains improved rates only depending on intrinsic dimension. Fu et al. [2024] studies conditional diffusion models for Hölder densities and Hu et al. [2024] further extends the framework to more advanced neural network architectures, e.g., diffusion transformers. Wibisono et al. [2024] establishes a minimax optimal rate to estimate Lipschitz score by kernel methods. With an $L^{2}$ accurate score estimator, several works provide the convergence rate of discrete samplers for diffusion models [Chen et al., 2022b, 2023a; Lee et al., 2023; Chen et al., 2024]. Combining score matching error and convergence of samplers, one can obtain an end-to-end distribution estimation error bound.
@@
The remarkable empirical success of transfer learning, meta-learning, and multi-task learning across a wide range of machine learning applications has been accompanied by gradual progress in their theoretical foundations, especially from the perspective of representation learning. To the best of our knowledge, Baxter [2000] is the first theoretical work on meta-learning. It assumes a universal environment to generate tasks with some shared features. Following this setting, Maurer et al. [2016] provides sample complexity bound for general supervised learning problem and Aliakbarpour et al. [2024] studies very few samples per task regime. Another line of research replaces the environment assumption and instead establishes connections between source tasks and target tasks through various notions of task diversity [Tripuraneni et al., 2020; Du et al., 2020; Tripuraneni et al., 2021; Watkins et al., 2023; Chua et al., 2021]. However, theoretical understandings of transfer learning for unsupervised learning are much more limited.
@@
Adapting pre-trained conditional diffusion models to specific tasks with limited data remains a challenge in varied application scenarios. Few-shot fine-tuning aims to bridge this gap by leveraging various techniques to adapt those models to a novel task with minimal data requirements [Ruiz et al., 2023; Giannone et al., 2022]. A promising paradigm is to use transfer (meta) learning by constructing a representation for conditions in all the tasks, which has been widely applied in image generation [Rombach et al., 2022; Ramesh et al., 2022; Sinha et al., 2021], reinforcement learning [He et al., 2023; Ni et al., 2023], inverse problem [Tewari et al., 2023; Chung et al., 2023], etc. Another recent work Yang et al. [2024] is closely related to this paper, proving that few-shot diffusion models can escape the curse of dimensionality by fine-tuning a linear encoder.
@@ 2 Preliminaries and Problem Setup
[…]
@@ 2.1 Conditional Diffusion Models
where $\{\overline{W}_{t}\}_{0\leq t\leq T}$ is a time-reversed Wiener process. Unfortunately, we don’t have access to the exact conditional score function $\nabla\log p_{T-t}$ and need to estimate it through neural networks. For any $(x,y)\sim\mathbb{P}$ and score estimator $s$, define the individual denoising score matching objective [Vincent, 2011] as
@@
Here $T_{0}>0$ is the early-stopping time. And the distribution of $\widehat{X}_{T-T_{0}}^{\leftarrow}$ is written as ^$\widehat{\mathbb{P}}(\cdot|y)$.
Note that we don’t apply the commonly used classifier-free guidance [Ho and Salimans, 2022] which has a tunable guidance strength since we mainly concentrate on sampling from conditional distribution instead of optimizing other objectives.
@@ 2.2 Transfer Diffusion Models via Learning Representation
Similar to Tripuraneni et al. [2020], our transfer learning procedures consist of two phases. In the pre-training phase, the goal is to learn a representation map $h_{*}$ through $nK$ samples from $K$ source distributions. Then during the fine-tuning phase, we learn the target distribution via $m$ new samples and the representation map learned in the pre-training phase.
@@
In the meta-learning setting, we further assume that all the distributions $\{\mathbb{P}^{k}\}_{k}$ are i.i.d. sampled from a meta distribution $\mathbb{P}_{\text{meta}}$. Here $\mathbb{P}_{\text{meta}}$ can be interpreted as a universal environment [Baxter, 2000; Maurer et al., 2016]. In this case, we posit the existence of a shared representation map that holds for all $\mathbb{P}\sim\mathbb{P}_{\text{meta}}$. And the performance benchmark is then defined as the expected error on the target distribution $\mathbb{P}^{0}\sim\mathbb{P}_{\text{meta}}$.
@@ 2.3 Deep ReLU Neural Network Family
We use feedforward neural networks to approximate the score function and representation map. Let $\sigma(x):=\max\{x,0\}$ be the ReLU activation. Define the neural network family
$NN_{f}(L,W,M,S,B,R,\gamma):=$$\Bigg{\{}f(x,w,t)=(A_{L}\sigma(\cdot)+b_{L})\circ\cdots\circ(A_{1}[x^{\top},w^{\top},t]^{\top}+b_{1}):$ (2.8)
$\quad A_{i}\in\mathbb{R}^{d_{i}\times d_{i+1}},b_{i}\in\mathbb{R}^{d_{i+1}},d_{L+1}=d_{x},\max d_{i}\leq W,\|f\|_{L^{\infty}}\leq M,$
$\quad\sum_{i=1}^{L}(\|A_{i}\|_{0}+\|b_{i}\|_{0})\leq S,\max\|A_{i}\|_{\infty}\vee\|b_{i}\|_{\infty}\leq B,$
$\quad\|f(x,w,t)-f(x,w^{\prime},t)\|\leq\gamma\|w-w^{\prime}\|_{\infty},\forall\ \|x\|_{\infty}\leq R,t\leq T\Bigg{\}},$
$NN_{h}(L,W,S,B):=$$\Bigg{\{}h(y)=(A_{L}\sigma(\cdot)+b_{L})\circ\cdots\circ(A_{1}y+b_{1}):A_{i}\in\mathbb{R}^{d_{i}\times d_{i+1}},$ (2.9)
$\quad b_{i}\in\mathbb{R}^{d_{i+1}},d_{L+1}=d_{y},\max d_{i}\leq W,\|h\|_{L^{\infty}([0,1]^{D_{y}})}\leq 1,$
$\quad\sum_{i=1}^{L}(\|A_{i}\|_{0}+\|b_{i}\|_{0})\leq S,\max\|A_{i}\|_{\infty}\vee\|b_{i}\|_{\infty}\leq B\Bigg{\}}.$
Throughout this paper, we let $\mathcal{F}^{0}=\mathcal{F}=NN_{f}(L_{f},W_{f},M_{f},S_{f},B_{f},R_{f},\gamma_{f})$ and $\mathcal{H}=NN_{h}(L_{h},W_{h},S_{h},B_{h})$ unless otherwise specified.
@@
In practice, $\mathcal{F}^{0}\subseteq\mathcal{F}$ may (and typically will) depend on $\widehat{{\bm{f}}}$ for parameter efficient fine-tuning (PEFT), e.g., LoRA [Hu et al., 2021]. This will substantially reduce the complexity of $\mathcal{F}^{0}$ and further improve sample efficiency. The analysis of PEFT is beyond the scope of this paper.
@@ 3 Statistical Guarantees for Transferring Score Matching Error
In this section, we present our main theoretical results, a statistical theory of transferring conditional score matching loss. We provide two upper bounds of score matching loss on target distribution, based on whether task diversity [Tripuraneni et al., 2020] is explicitly assumed. Throughout this paper, we make the following standard and mild regularity assumptions on the initial data distribution $\mathbb{P}$ and the representation map $h_{*}$.
@@
For any source and target distribution $\mathbb{P}$, $\mathbb{P}$ is supported on $\mathbb{R}^{d_{x}}\times[0,1]^{D_{y}}$ and admits a continuous density $p(x,y)\in\mathcal{C}^{2}(\mathbb{R}^{d_{x}}\times[0,1]^{D_{y}})$. Moreover, the conditional distribution $p(x|y)\leq C_{1}\exp(-C_{2}\|x\|^{2})$ for some constant $C_{1},C_{2}$.

### v2

@@ Abstract
While conditional diffusion models have achieved remarkable success in various applications, they require abundant data to train from scratch, which is often infeasible in practice. To address this issue, transfer learning has emerged as an essential paradigm in small data regimes. Despite its empirical success, the theoretical underpinnings of transfer learning conditional diffusion models remain unexplored. In this paper, we take the first step towards understanding the sample efficiency of transfer learning conditional diffusion models through the lens of representation learning. Inspired by practical training procedures, we assume that there exists a low-dimensional representation of conditions shared across all tasks. Our analysis shows that with a well-learned representation from source tasks, the sample complexity of target tasks can be reduced substantially. Numerical experiments are also conducted to verify our results.
@@ 1 Introduction
Conditional diffusion models (CDMs) utilize a user-defined condition to guide the generative process of diffusion models (DMs) to sample from the desired conditional distribution. In recent years, CDMs have achieved groundbreaking success in various generative tasks, including text-to-image generation (Ho et al., 2020; Song et al., 2020; Ho and Salimans, 2022; Rombach et al., 2022), reinforcement learning (Janner et al., 2022; Chi et al., 2023; Wang et al., 2022; Reuss et al., 2023), time series (Tashiro et al., 2021; Rasul et al., 2021), and life science (Song et al., 2021; Watson et al., 2022; Gruver et al., 2024; Guo et al., 2024).
Training a CDM from scratch requires a large amount of data to achieve good generalization. However, in practical scenarios, users often have access to only limited data for the target distribution due to cost or risk concerns, making the model prone to over-fitting. In such small data regime, transfer learning has emerged as a predominant paradigm (Moon et al., 2022; Ruiz et al., 2023; Xie et al., 2023; Han et al., 2023). By leveraging knowledge acquired during pre-training on large source datasets, transfer learning enhances the performance of fine-tuning on target tasks, facilitating few-shot learning and significantly improving practicality.
@@
Text-to-Image (Esser et al., 2024) 2-8B 4.7B
Text-to-Audio (Liu et al., 2024) 350-750M 750M
Robotic Control (Chi et al., 2023) 9M 20-45M
@@
Among the successful applications of transfer learning CDMs, the conditions are typically high-dimensional vectors with embedded low-dimensional representations (features) that encapsulate all the information required for inference. In addition, these representations are likely to be task-agnostic, enabling effective knowledge transfer. For example, in text-to-image generation, the text input is inherently in high-dimensional space, but contains low-dimensional semantic information such as object attributes, spatial relationships, despite the differences of styles or contents in different image distributions. To take advantage of this structure, condition encoders are often frozen in the fine-tuning stage (Rombach et al., 2022; Esser et al., 2024), which typically constitutes a significant portion of the overall model (see Table 1).
@@
There are some recent works attempting to study the theoretical underpinnings of CDMs (Fu et al., 2024; Jiao et al., 2024; Hu et al., 2024), but focus on single task training. Notably, Yang et al. (2024) investigates transfer learning DMs under the assumption that the data is a linear transformation of a low-dimensional latent variable following the same distribution across all tasks. However, fine-tuning merely the data encoder is not a widely adopted training approach in practice.
@@
  • In Section 3, we establish the first generalization guarantee for transferring score matching error in CDMs, showing that transfer learning can reduce the sample complexity for learning condition encoder in the target task. This is aligned with existing transfer learning theory in supervised learning. Specifically, we present two results in Theorem 3.4 and Theorem 3.6, under the settings of task diversity assumption and meta-learning [^1] In practice, the terms such as transfer learning, meta-learning, learning-to-learn, etc., often refer to the same training paradigm, i.e., to fine-tune on target tasks with limited data using knowledge from source tasks. However, in the theoretical framework, we use meta-learning to emphasize that target tasks and source tasks are randomly sampled from a meta distribution (Baxter, 2000), whereas in transfer learning, the tasks are fixed., respectively. On the technical side, we develop a novel approach to tackle Lipschitz continuity under weaker assumptions on data distribution in Lemma 3.1, which may be of independent interest for the analysis of even single-task diffusion models.
@@
In Section A, we further utilize our results to establish statistical guarantees in practical applications of CDMs. In particular, we investigate amortized variational inference (Theorem A.1) and behavior cloning (Theorem A.2), and present guarantees in terms of posterior estimation and optimality gap, laying the theoretical foundations of transfer learning CDMs in practice. We also conduct numerical experiments in Section 5 to verify our results.
@@ 1.1 Related Works
Recently, some works analyze the score approximation theory via deep neural networks and corresponding sample complexity bounds for diffusion models. Oko et al. (2023) considers distributions with density in Besov space and supported on bounded domain. Chen et al. (2023b) assumes the data distribution lies in a low-dimensional linear subspace and obtains improved rates only depending on intrinsic dimension. Fu et al. (2024) studies conditional diffusion models for Hölder densities and Hu et al. (2024) further extends the framework to more advanced neural network architectures, e.g., diffusion transformers. Wibisono et al. (2024) establishes a minimax optimal rate to estimate Lipschitz score by kernel methods. With an $L^{2}$ accurate score estimator, several works provide the convergence rate of discrete samplers for diffusion models (Chen et al., 2022b, 2023a; Lee et al., 2023; Chen et al., 2024). Combining score matching error and convergence of samplers, one can obtain an end-to-end distribution estimation error bound.
@@
The remarkable empirical success of transfer learning, meta-learning, and multi-task learning across a wide range of machine learning applications has been accompanied by gradual progress in their theoretical foundations, especially from the perspective of representation learning. To the best of our knowledge, Baxter (2000) is the first theoretical work on meta-learning. It assumes a universal environment to generate tasks with some shared features. Following this setting, Maurer et al. (2016) provides sample complexity bound for general supervised learning problem and Aliakbarpour et al. (2024) studies very few samples per task regime. Another line of research replaces the environment assumption and instead establishes connections between source tasks and target tasks through various notions of task diversity (Tripuraneni et al., 2020; Du et al., 2020; Tripuraneni et al., 2021; Watkins et al., 2023; Chua et al., 2021). However, theoretical understandings of transfer learning for unsupervised learning are much more limited.
@@
Adapting pre-trained conditional diffusion models to specific tasks with limited data remains a challenge in varied application scenarios. Few-shot fine-tuning aims to bridge this gap by leveraging various techniques to adapt those models to a novel task with minimal data requirements (Ruiz et al., 2023; Giannone et al., 2022). A promising paradigm is to use transfer (meta) learning by constructing a representation for conditions in all the tasks, which has been widely applied in image generation (Rombach et al., 2022; Ramesh et al., 2022; Sinha et al., 2021), reinforcement learning (He et al., 2023; Ni et al., 2023), inverse problem (Tewari et al., 2023; Chung et al., 2023), etc. Another work Yang et al. (2024) is closely related to this paper, proving that few-shot diffusion models can escape the curse of dimensionality by fine-tuning a linear encoder.
@@ 2 Preliminaries and Problem Setup
We use $x$ and $y$ to denote the data and conditions, respectively. The blackboard bold letter $\mathbb{P}$ represents the joint distribution of $(x,y)$, while the lowercase $p$ denotes its density function. The superscript $k$ indicates the task index, and the subscript $i$ means the sample index. The norm $\|\cdot\|$ refers to the $\ell_{2}$-norm for vectors and the spectral norm for matrices. For the hypothesis class $\mathcal{F}$, we use $\mathcal{F}^{\otimes K}$ to refer its $K$-fold Cartesian product. For any $a,b\in\mathbb{R}$, $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$. Finally, we use standard $\mathcal{O}(\cdot),\Omega(\cdot)$ to omit constant factors.
@@ 2.1 Conditional Diffusion Models
where $\{\overline{W}_{t}\}_{0\leq t\leq T}$ is a time-reversed Wiener process. Unfortunately, we don’t have access to the exact conditional score function $\nabla\log p_{T-t}$ and need to estimate it through neural networks. For any $(x,y)\sim\mathbb{P}$ and score estimator $s$, define the individual denoising score matching objective (Vincent, 2011) as
@@
Here $T_{0}>0$ is the early-stopping time. And the distribution of $\widehat{X}_{T-T_{0}}^{\leftarrow}$ is written as $\widehat{\mathbb{P}}(\cdot|y)$.
Note that we don’t apply the commonly used classifier-free guidance (Ho and Salimans, 2022) which has a tunable guidance strength since we mainly concentrate on sampling from conditional distribution instead of optimizing other objectives.
@@ 2.2 Transfer Diffusion Models via Learning Representation
Similar to Tripuraneni et al. (2020), our transfer learning procedures consist of two phases. In the pre-training phase, the goal is to learn a representation map $h_{*}$ through $nK$ samples from $K$ source distributions. Then during the fine-tuning phase, we learn the target distribution via $m$ new samples and the representation map learned in the pre-training phase.
@@
In the meta-learning setting, we further assume that all the distributions $\{\mathbb{P}^{k}\}_{k}$ are i.i.d. sampled from a meta distribution $\mathbb{P}_{\text{meta}}$. Here $\mathbb{P}_{\text{meta}}$ can be interpreted as a universal environment (Baxter, 2000; Maurer et al., 2016). In this case, we posit the existence of a shared representation map that holds for all $\mathbb{P}\sim\mathbb{P}_{\text{meta}}$. And the performance benchmark is then defined as the expected error on the target distribution $\mathbb{P}^{0}\sim\mathbb{P}_{\text{meta}}$.
@@ 2.3 Deep ReLU Neural Network Family
We use feedforward neural networks to approximate the score function and representation map. Let $\sigma(x):=\max\{x,0\}$ be the ReLU activation. Define the score network family $NN_{f}(L,W,M,S,B,R,\gamma):=\big\{f(x,w,t)=(A_{L}\sigma(\cdot)+b_{L})\circ\cdots\circ(A_{1}[x^{\top},w^{\top},t]^{\top}+b_{1}):A_{i}\in\mathbb{R}^{d_{i}\times d_{i+1}},b_{i}\in\mathbb{R}^{d_{i+1}},d_{L+1}=d_{x},\max d_{i}\leq W,\|f\|_{L^{\infty}}\leq M,\sum_{i=1}^{L}(\|A_{i}\|_{0}+\|b_{i}\|_{0})\leq S,\max\|A_{i}\|_{\infty}\vee\|b_{i}\|_{\infty}\leq B,\|f(x,w,t)-f(x,w^{\prime},t)\|\leq\gamma\|w-w^{\prime}\|_{\infty},\forall\ \|x\|_{\infty}\leq R,t\leq T\big\}$, and encoder network $NN_{h}(L,W,S,B):=\big\{h(y)=(A_{L}\sigma(\cdot)+b_{L})\circ\cdots\circ(A_{1}y+b_{1}):A_{i}\in\mathbb{R}^{d_{i}\times d_{i+1}},b_{i}\in\mathbb{R}^{d_{i+1}},d_{L+1}=d_{y},\max d_{i}\leq W,\|h\|_{L^{\infty}([0,1]^{D_{y}})}\leq 1,\sum_{i=1}^{L}(\|A_{i}\|_{0}+\|b_{i}\|_{0})\leq S,\max\|A_{i}\|_{\infty}\vee\|b_{i}\|_{\infty}\leq B\big\}.$ Throughout this paper, we let $\mathcal{F}^{0}=\mathcal{F}=NN_{f}(L_{f},W_{f},M_{f},S_{f},B_{f},R_{f},\gamma_{f})$ and $\mathcal{H}=NN_{h}(L_{h},W_{h},S_{h},B_{h})$ unless otherwise specified.
@@
In practice, $\mathcal{F}^{0}\subseteq\mathcal{F}$ may (and typically will) depend on $\widehat{{\bm{f}}}$ for parameter efficient fine-tuning (PEFT), e.g., LoRA (Hu et al., 2021). This will substantially reduce the complexity of $\mathcal{F}^{0}$ and further improve sample efficiency. The analysis of PEFT is beyond the scope of this paper.
@@ 3 Statistical Guarantees for Transferring Score Matching Error
In this section, we present our main theoretical results, a statistical theory of transferring the conditional score matching loss. We provide two upper bounds of the score matching loss on target distribution, based on whether task diversity (Tripuraneni et al., 2020) is explicitly assumed. Our analysis introduces novel techniques to address the smoothness properties of the noised data distribution—a challenge that remains nontrivial even in single-task settings. Additionally, we extend the classical theory of local Rademacher complexity to quantify the empirical estimation error.
Throughout this paper, we make the following standard and mild regularity assumptions (Tripuraneni et al., 2020; Chen et al., 2023b) on the initial data distribution $\mathbb{P}$ and the representation map $h_{*}$.
@@
For any source or target distribution $\mathbb{P}$, $\mathbb{P}$ is supported on $\mathbb{R}^{d_{x}}\times[0,1]^{D_{y}}$ and admits a continuous density $p(x,y)\in\mathcal{C}^{2}(\mathbb{R}^{d_{x}}\times[0,1]^{D_{y}})$. Moreover, the conditional distribution $p(x|y)\leq C_{1}\exp(-C_{2}\|x\|^{2})$ for some constant $C_{1},C_{2}$.